# Context-aware Fairness Evaluation and Mitigation in LLMs

### v1

@@ 1 Introduction
Traditional mitigation has centered on training-time interventions—supervised fine-tuning, RLHF, self-debiasing, and data filtering—which can be effective but are computationally costly, irreversible after deployment, and inflexible to new requirements, since they demand full retraining or heavy data curation Shirafuji et al. (2024); Bayasi et al. (2024); Tang et al. (2025); Schick et al. (2021); Ok et al. (2025); Gupta et al. (2025); Defrance et al. (2025). In contrast, inference-level methods—prompt engineering, sequential calibration, vector steering, neuron pruning, and other representation-level controls—are lightweight and reversible, yet as Figure 1 illustrates they typically act at shallow levels, stay static across turns, and are largely tailored to single-turn use Gupta et al. (2024); Fei et al. (2023); Chen et al. (2025); Yang et al. (2023); Li et al. (2025); Yi et al. (2025). Within this space, pruning-based techniques directly manipulate neuron activations and offer interpretability and efficiency—e.g., attribution-based pruning for task compression, structured neuron pruning for interpretable adaptation, and pruning for specialization Yang et al. (2024a); Ma et al. (2023); Dufort-Labbé et al. (2024). However, pruning remains globally applied and irreversible: importance scores are computed offline, risking over-pruning and semantic loss, and failing to capture the dynamic, cumulative nature of bias that emerges from contextual dependencies and dialogue history in multi-turn conversations Zhang et al. (2025).
@@
To assess this, we evaluate single and multi turn prompts. Single-turn evaluation and mitigation assesses a model’s immediate responses to isolated prompts, providing a controlled setting to analyse surface-level bias. This provides an understanding of a model’s baseline behavior and ensures fairness at the response level. For single-turn evaluation and mitigation, we focus on an ideological or representational bias in isolated prompts, offering a controlled way to assess and correct stance-level tendencies without conversational context. Political Compass Test (PCT) [^1] https://www.politicalcompass.org, are valuable for revealing how models frame political or moral positions, enabling calibration of surface-level bias at the response level. However, it remains limited in treating bias as a static property of a single output rather than a dynamic process that evolves through context, memory, and user interaction. In real conversational settings, bias often accumulates or re-emerges across turns as earlier activations influence subsequent responses. This is the need for multi-turn evaluation and mitigation, which captures how bias propagates, compounds, or shifts over dialogue history, providing a more realistic and comprehensive view of fairness in conversational LLMs. To achieve this task, we focus on stereotype and toxicity bias, as they capture complementary aspects of social bias. Stereotypes reflect implicit associations, while toxicity measures explicit, harmful content. Combined, they cover the most influential forms of bias affecting fairness and safety in conversational LLMs across six demographic attributes: Gender, Race, Appearance, Disability, Religion, and Age in multi-turn dialogue settings [^2] https://github.com/FanZT6/FairMT-bench for more details in Section 4.2. The research contributions are:
  • We formalize bias as a dynamic, accumulating phenomenon, separating local vs. memory-borne bias and introducing diagnostics for carry-over across dialogue turns to ensure turn-aware fairness.
  • An inference-time, reversible framework that detects, probes, and selectively gates bias-carrying activations as generation unfolds—no retraining and hard neuron suppression or pruning.
  • Multilingual, multi-turn wins with utility intact. On a Multilingual, multi-turn dataset, dynamic neuron masking consistently reduces bias while preserving fluency, faithfulness, and relevance
This positions a new paradigm for inference-level fairness in conversational LLMs.
@@ 3 Methodology
Behavioral detection is the starting point for identifying when biased or harmful behavior appears in a model’s response. It provides a clear signal of bias at the output level, helping connect what the model says to the internal activations that caused it. We first estimate a turn-level bias score $S_{t}$ Dai et al. (2025), using LLM-based evaluation using GPT-3.5Turbo as the main assessor with Claude as an auxiliary judge Fan et al. (2024). We combined the bias scores as a weighted average to ensure reliability. We perform human annotation on 20% of samples to verify model judgments and audit disagreements. We employed two annotators for 20% of the data and measured reliability using Cohens’ K. Agreement showed substantial agreement.
@@
Table 1: Political inclination on PCT Across Languages and Models
3.2 Bias Neuron Identification
Bias neuron identification finds the neurons responsible for biased behavior to know the model to address the root cause for bias emergence, we attribute $S_{t}$ to specific neurons using integrated gradients Yang et al. (2024b). We compute an importance score for each neuron that is decomposed into two complementary components: $m_{\text{local}}$, capturing neuron relevance within the current turn, and $m_{\text{carry}}$, quantifying carry-over effects from earlier dialogue turns. This decomposition enables us to distinguish transient, turn-specific activations from memory-driven bias propagation across the conversation.
3.3 Aggregation of Bias Scores
[…]
Token aggregation.
Because activations are token-local, we pool over tokens to obtain a single score for $(\iota,x_{j})$:
Instance aggregation.
Tasks contain multiple instances $D=\{(x_{j},y_{j})\}_{j=1}^{N}$. We weight instances by a confusion score $\alpha(\iota,x_{j})$, scaled by the dialogue-conditioned importance from $m_{\text{local}}$ and $m_{\text{carry}}$, such that persistent carry-over activations are not diluted. The aggregated score is
with normalized weights
where $\alpha(\iota,x_{j})=\Pr(\hat{y}_{j}\mid\iota,x_{j})$.
Instruction aggregation.
To be robust across paraphrases, we average over a set of semantically equivalent instructions $I$ (with $|I|=M$):
$B^{(I,D)}_{i}(h)\;=\;\frac{1}{M}\sum_{\iota\in I}B^{(\iota,D)}_{i}(h).$ (4)
3.4 Concept-Based Tesing
Figure 4: Bias progression across dialogue turns and tasks. (Left) Turn-wise comparison of bias across different models shows an increase in later turns. (Middle) Bias across demographic groups highlights stronger signals for religion, race, and gender in multi-turn contexts.(Right) model–task interactions indicate task-specific vulnerability, with Scattered Questions (SQ) and Inference Misinformation (IM) producing higher bias scores across several models.
3.5 Dynamic Neuron Masking
Instead of permanently removing neurons, this step adaptively adjusts their influence during generation because bias in conversations changes with context and memory so Dynamic neuron masking is needed to preserve the model’s fairness and consistency. Finally, we apply controlled masking that adaptively gates neurons during decoding. For each layer $l$, the gating coefficient is $g_{l}^{(t)}=\sigma(\alpha S_{t}+\beta C_{t})$, where $\sigma$ is the sigmoid function and $\alpha,\beta$ are tunable weights. This mechanism dynamically suppresses high-bias neurons while retaining information essential for coherence and fluency. The proposed Dynamic Neuron Masking framework requires no model retraining and adds only a small per-token cost for scoring and gating, comparable to other real-time steering methods Chen et al. (2025); Gupta et al. (2025).
The Bias Score is defined as $S_{t}=\dfrac{n_{b}}{n_{t}}$, where $n_{b}$ is the number of biased texts and $n_{t}$ is the total number of texts. The Toxicity metric is $T=\dfrac{n_{tox}}{n_{t}}$, where $n_{tox}$ denotes the number of toxic texts. Knowledge Retention is given by $C_{t}=\dfrac{n_{kr}}{n_{t}}$, where $n_{kr}$ is the number of texts without knowledge attrition. Faithfulness is measured as $F=\dfrac{n_{truth}}{n_{claims}}$, with $n_{truth}$ the number of truthful claims and $n_{claims}$ the total number of claims. Finally, Answer Relevancy is $R=\dfrac{n_{rel}}{n_{stmts}}$, where $n_{rel}$ is the number of relevant statements and $n_{stmts}$ the total number of statements.
Together, these metrics provide a balanced protocol: $S_{t}$ and $T$ quantify fairness and safety, $C_{t}$ measures factual memory retention, and $F$ and $R$ ensure responses remain contextually grounded and aligned with user intent to ensure coherence and relevance DeepEval (2025).
4 Experiments
4.1 Baseline Comparison
@@
Comparisons include comprehensive inference-time baselines based on Prompt Engineering Chisca et al. (2024), logit filtering Fei et al. (2023), vector steering Siddique et al. (2025), and neuron pruning Yang et al. (2024a), results shown in Table 2.
4.2 Datasets
We evaluate our framework on two complementary benchmarks to capture both single-turn and multi-turn conversational bias.
Static: Single Turn
The multilingual PCT dataset Nadeem et al. (2025a)contains 62 ideological statements translated into five low-resource languages (Urdu, Punjabi, Sindhi, Pashto, Balochi) plus English Thapa et al. (2023); Helwe et al. (2025). Each item is a single-turn prompt eliciting an agree/disagree style response, enabling controlled measurement of political-ideology bias across languages of Pakistan Rahman (1996, 2011); Umrani and Bughio (2020); Abbas and Bidin (2022).
Dynamic: Multi Turn
@@ 4.2 Datasets
FairMT-Bench Fan et al. (2024) extends bias evaluation to multi-turn dialogue. It comprises 10K conversations across fairness critical categories such as anaphora ellipsis, scatter questions, jailbreak tips, inference misinformation, negative feedback, and fixed format. Bias accumulation and memory carry-over are explicitly annotated, making it ideal for testing context-aware mitigation. Automatic evaluation uses an ensemble: GPT-3.5-Turbo as the primary judge, Llama Guard as the classifier Fan et al. (2024), and Claude as an auxiliary validator, see model details in Appendix A.1.
4.3 Bias Exists
@@
Figure 5: Comparison of effective bias scores across dialogue turns, baseline (red, no masking) and mitigation (green, with masking) showing higher bias accumulation in the baseline (no masking) and substantial reduction with masking applied.
PCT-SingleTurn
Model Method English Urdu Punjabi Sindhi Balochi Pashto
@@
FairMT-MultiTurn
Model Method AE SQ JT IM NF $F$
@@
Table 2: Comparison of bias mitigation methods and percentage of bias score (lower is better). Values in parentheses show 95% bootstrap confidence intervals. PCT-SingleTurn results are across six languages, and FairMT-MultiTurn across six tasks.

### v2

@@ 1 Introduction
Traditional mitigation has centered on training-time interventions supervised fine-tuning, RLHF, self debiasing, and data filtering which can be effective but are computationally costly, irreversible after deployment, and inflexible to new requirements, since they demand full retraining or heavy data curation Shirafuji et al. (2024); Bayasi et al. (2024); Tang et al. (2025); Schick et al. (2021); Ok et al. (2025); Gupta et al. (2025); Defrance et al. (2025). In contrast, inference-level methods prompt engineering, sequential calibration, vector steering, neuron pruning, and other representation level controls are lightweight and reversible, yet as Figure 1 illustrates they typically act at shallow levels, stay static across turns, and are largely tailored to single-turn use Gupta et al. (2024); Fei et al. (2023); Chen et al. (2025); Yang et al. (2023); Li et al. (2025); Yi et al. (2025). Within this space, pruning-based techniques directly manipulate neuron activations and offer interpretability and efficiency e.g., attribution-based pruning for task compression, structured neuron pruning for interpretable adaptation, and pruning for specialization Yang et al. (2024a); Ma et al. (2023); Dufort-Labbé et al. (2024). However, pruning remains globally applied and irreversible: importance scores are computed offline, risking over-pruning and semantic loss, and failing to capture the dynamic, cumulative nature of bias that emerges from contextual dependencies and dialogue history in multi-turn conversations Zhang et al. (2025). A key limitation of existing pruning-based mitigation methods is the assumption that neurons associated with biased outputs are inherently undesirable and should be permanently removed. In practice, however, neuron activations in large language models are context dependent and multifunctional; the same neuron that amplifies biased content under certain dialogue histories may encode legitimate semantic, factual, or discourse information in neutral or corrective contexts. In multi-turn conversations, bias often emerges gradually through priming, anaphora, and memory carry-over, rather than being uniformly present at every turn. As a result, static pruning over-suppresses neurons even when bias is absent, leading to semantic loss, reduced answer relevance, and degraded coherence. This effect is particularly pronounced in long-horizon dialogue, where earlier activations are reused and repurposed across turns. These observations motivate a dynamic, memory-aware intervention that conditions neuron suppression on when biased behavior is observed, rather than treating bias as a fixed neuron property.
@@
To assess this, we evaluate fairness as a core pillar of alignment LLMs, requiring the measurement and mitigation of cultural, ideological, demographic, and political biases to guide models toward culturally grounded neutrality. This entails reducing ideological bias, political inclination, and harmful stereotypes without compromising factual knowledge, coherence, or local relevance challenges that are mainly pronounced in multilingual, multi-turn dialogue, where bias is context-dependent and static safeguards often fail. We evaluate state-of-the-art LLMs across multilingualism , using Political Compass Test (PCT) [^1] https://www.politicalcompass.org and BBQ [^2] https://github.com/nyu-mll/BBQ/tree/main/data. Then, we apply existing bias mitigation methods at the representation level, enabling effective single-turn bias reduction. However, in multi-turn conversations, bias re-emerges through dialogue history and repeated steering degrades coherence and fluency. However, it remains limited in treating bias as a static property of a single-turn output rather than a dynamic process that evolves through context, memory, and user interaction. In real conversational settings, bias often accumulates or re-emerges across turns as earlier activations influence subsequent responses. This is the need for multi-turn evaluation and mitigation, which captures how bias propagates, compounds, or shifts over dialogue history, providing a more realistic and comprehensive view of fairness in conversational LLMs. To achieve this task, we focus on stereotype and toxicity bias, as they capture complementary aspects of social bias. Stereotypes reflect implicit associations, while toxicity measures explicit, harmful content. Combined, they cover the most influential forms of bias affecting fairness and safety in conversational LLMs across six demographic attributes: Gender, Race, Appearance, Disability, Religion, and Age in multi-turn dialogue settings [^3] https://github.com/FanZT6/FairMT-bench for more details in Section 4. The contributions are: We formalize bias as a dynamic, accumulating phenomenon, separating local vs. memory-borne bias and introducing diagnostics for carry-over across dialogue turns to ensure turn-aware fairness.
An inference-time, reversible framework that detects, probes, and selectively gates bias-carrying activations as generation unfolds, no retraining and hard neuron suppression or pruning.
Multilingual, multi-turn wins with utility intact. On a Multilingual, multi-turn dataset, dynamic neuron masking consistently reduces bias while preserving fluency, faithfulness, and relevance This positions a new paradigm for inference-level fairness in conversational LLMs.
@@ 3 Methodology
Behavioral detection is the starting point for identifying when biased or harmful behavior appears in a model’s response. It provides a clear signal of bias at the output level, helping connect what the model says to the internal activations that caused it. We first estimate a turn-level bias score $S_{t}$ Dai et al. (2025), using LLM-based evaluation GPT-3.5Turbo as the main assessor with Claude as an auxiliary judge Fan et al. (2024). We combined the bias scores as a weighted average to ensure reliability. We perform human annotation on 20% of samples to verify model judgments and audit disagreements. We employed two annotators for 20% of the data and measured reliability using Cohens’ K 0.71 (95% CI: [0.55, 0.83]) for Multi-turn and K 0.98 (95% CI: [0.95, 1.00]) for multilingual PCT dataset. Agreement showed substantial agreement, for more details see Appendix 10.1.
3.2 Bias Neuron Identification
To identify neurons responsible for biased behavior, we attribute the turn-level bias score $S_{t}$ to internal neuron activations using integrated gradients Yang et al. (2024b). For each neuron, we compute an importance score that measures its contribution to biased outputs by aggregating attribution strength across tokens and instances. This score is decomposed into two components: $m_{\text{local}}$, capturing the neuron’s relevance in the current turn, and $m_{\text{carry}}$, capturing persistent influence from earlier dialogue turns. This decomposition distinguishes transient, context-specific bias from memory-driven bias propagation, enabling targeted and context-aware mitigation.
3.3 Aggregation of Bias Scores
[…]
Instance aggregation: Tasks contain multiple instances $D=\{(x_{j},y_{j})\}_{j=1}^{N}$. We weight instances by a confusion score $\alpha(\iota,x_{j})$, scaled by the dialogue-conditioned importance from $m_{\text{local}}$ and $m_{\text{carry}}$, such that persistent carry-over activations are not diluted. The aggregated score with normalized weights are:
where $\alpha(\iota,x_{j})=\Pr(\hat{y}_{j}\mid\iota,x_{j})$.
Instruction aggregation: To be robust across paraphrases, we average over a set of semantically equivalent instructions $I$ (with $|I|=M$):
$B^{(I,D)}_{i}(h)\;=\;\frac{1}{M}\sum_{\iota\in I}B^{(\iota,D)}_{i}(h).$ (4)
3.4 Concept-Based Testing
Figure 4: Evaluation of Bias progression across dialogue turns and tasks. (Left) Turn-wise comparison of bias across different models shows an increase in later turns. (Middle) Bias across demographic groups highlights stronger signals for religion, race, and gender in multi-turn contexts.(Right) model task interactions indicate task-specific vulnerability, with Scattered Questions (SQ) and Inference Misinformation (IM) producing higher bias scores across several models.
3.5 Dynamic Neuron Masking
Instead of permanently removing neurons, this step adaptively adjusts their influence during generation because bias in conversations changes with context and memory so Dynamic neuron masking is needed to preserve the model’s fairness and consistency, more details in Appendix 10.3. Finally, we apply controlled masking that adaptively gates neurons during decoding. For each layer $l$, the gating coefficient is $g_{l}^{(t)}=\sigma(\alpha S_{t}+\beta C_{t})$, where $\sigma$ is the sigmoid function and $\alpha,\beta$ are tunable weights. This mechanism dynamically suppresses high-bias neurons while retaining information essential for coherence and fluency. The proposed Dynamic Neuron Masking framework requires no model retraining and adds only a small per-token cost for scoring and gating, comparable to other real-time steering methods Chen et al. (2025); Gupta et al. (2025).
@@ 3.5 Dynamic Neuron Masking
Table 1: Evaluation of political inclination on PCT across Languages and Models, value represent stance score on social and economic axis.
4 Experiment Settings
Baseline Comparison:
@@ 4 Experiment Settings
Comparisons include comprehensive inference-time baselines based on Prompt Engineering Chisca et al. (2024), logit filtering Fei et al. (2023), vector steering Siddique et al. (2025), and neuron pruning Yang et al. (2024a), results in Table 2. For evaluation we used Mistral-7B-Instruct-v0.2-7B and DeepSeek-Chat-7B, see more details in Appendix 10.4.
Datasets:
@@
We evaluate our framework on two complementary benchmarks to capture both single-turn and multi-turn conversational bias. Static and Single Turn: The multilingual PCT dataset Nadeem et al. (2025a)contains 62 ideological statements translated into five low-resource languages (Urdu, Punjabi, Sindhi, Pashto, Balochi) plus English Thapa et al. (2023); Helwe et al. (2025). Each item is a single-turn prompt eliciting an agree/disagree style response, enabling controlled measurement of political-ideology bias across languages of Pakistan Rahman (1996, 2011); Umrani and Bughio (2020); Abbas and Bidin (2022). Another benchmark BBQ evaluates social bias in single-turn multiple-choice reasoning across six protected attributes Parrish et al. (2022). Dynamic and Multi Turn: FairMT-Bench Fan et al. (2024) extends bias evaluation to multi-turn dialogue. It comprises 10K conversations across fairness critical categories such as anaphora ellipsis, scatter questions, jailbreak tips, inference misinformation, negative feedback, and fixed format. Bias accumulation and memory carry-over are explicitly annotated, making it ideal for testing context-aware mitigation. Automatic evaluation uses an ensemble: GPT-3.5-Turbo as the primary judge, Llama Guard as the classifier Fan et al. (2024), and Claude as an auxiliary validator, see model details in Appendix 10.4.
Evaluation Metrics:
To evaluate the performance the metrics for Bias Score is defined as $S_{t}=\dfrac{n_{b}}{n_{t}}$, where $n_{b}$ is the number of biased texts and $n_{t}$ is the total number of texts. The Toxicity metric is $T=\dfrac{n_{tox}}{n_{t}}$, where $n_{tox}$ denotes the number of toxic texts. Knowledge Retention is given by $K_{t}=\dfrac{n_{kr}}{n_{t}}$, where $n_{kr}$ is the number of texts without knowledge attrition. Faithfulness is measured as $F=\dfrac{n_{truth}}{n_{claims}}$, with $n_{truth}$ the number of truthful claims and $n_{claims}$ the total number of claims. Finally, Answer Relevancy is $R=\dfrac{n_{rel}}{n_{stmts}}$, where $n_{rel}$ is the number of relevant statements and $n_{stmts}$ the total number of statements. Together, these metrics provide a balanced protocol: $S_{t}$ and $T$ quantify fairness and safety, $C_{t}$ measures factual memory retention, and $F$,$K$ and $R$ ensure responses remain contextually grounded and aligned with user intent to ensure coherence and relevance DeepEval (2025).
Bias Exists:
@@
Figure 5: Mitigation comparison of effective bias scores across dialogue turns, baseline (red, no masking) and mitigation (green, with masking) showing higher bias accumulation in the baseline (no masking) and substantial reduction with masking applied.
Model Method evaluated for PCT-SingleTurn English Urdu Punjabi Sindhi Balochi Pashto
@@
Model Method evaluated for BBQ-SingleTurn Age Disability Race Appearance Gender Religion
Mistral Baseline 34.8 (±1.0) 27.6 (±2.3) 45.9 (±2.6) 33.2 (±2.9) 42.8 (±2.4) 53.1 (±3.1)
Prompt Engineering 34.5 (±1.8) 32.9 (±1.9) 38.7 (±2.3) 38.4 (±2.7) 29.6 (±2.5) 25.8 (±2.6)
Output Filtering 31.7 (±1.6) 23.2 (±1.7) 32.1 (±2.1) 24.6 (±1.5) 36.8 (±2.4) 41.9 (±2.4)
Steering Vectors Ensembles 28.9 (±1.5) 21.0 (±1.6) 19.8 (±1.0) 38.2 (±1.2) 34.1 (±1.0) 38.6 (±2.3)
Neuron Pruning 25.4 (±1.3) 27.9 (±1.5) 35.6 (±1.9) 24.3 (±2.1) 21.5 (±2.9) 32.7 (±2.2)
Dynamic Neurons Masking (Proposed) 19.6 (±1.1) 22.8 (±1.3) 28.7 (±1.6) 20.9 (±1.5) 26.4 (±1.4) 21.2 (±1.6)
DeepSeek Baseline 26.5 (±1.9) 33.8 (±2.1) 39.2 (±2.4) 36.8 (±2.3) 40.9 (±2.6) 33.7 (±2.7)
Prompt Engineering 21.2 (±1.7) 24.5 (±1.8) 35.0 (±1.1) 23.6 (±1.0) 36.2 (±2.3) 28.4 (±2.4)
Output Filtering 29.4 (±1.6) 32.1 (±1.7) 21.7 (±2.0) 41.9 (±2.0) 30.8 (±2.1) 32.6 (±2.2)
Steering Vectors Ensembles 25.8 (±1.5) 27.4 (±1.4) 25.1 (±1.8) 39.2 (±1.9) 45.7 (±2.0) 48.9 (±2.2)
Neuron Pruning 22.1 (±1.3) 24.8 (±1.5) 31.9 (±1.7) 36.1 (±1.8) 42.6 (±1.9) 44.5 (±2.1)
Dynamic Neurons Masking (Proposed) 16.3 (±1.0) 18.9 (±1.1) 19.7 (±1.3) 27.6 (±1.4) 24.1 (±1.3) 21.8 (±1.4)
Model Method evaluated for FairMT-Bench Anaphora Elipsis Scattered Questions Jailbreak Tips Inte. Misinfo Negative Feedback Fixed Format
@@
Table 2: Comparison of bias mitigation methods and percentage of bias score (lower is better). Values in parentheses denote 95% bootstrap confidence intervals. The baseline exhibits consistently higher bias due to the absence of representation-level intervention, where as other methods reduce bias as its static or permanent interventions which can degrade coherence and utility due to over-suppression of shared representations. In contrast, the proposed method significant bias reduction by dynamically suppressing bias-carrying neuron activations in a context- and memory-aware manner, preventing bias re-emergence while preserving coherence and utility.
5 Result and Analysis
In the Table 2, the results show without any intervention, both Mistral and DeepSeek carry a high level of bias across languages and dialogue tasks. Simple fixes like prompt engineering or output filtering help a little but don’t fully solve the problem, while methods such as steering vectors ensembles, and pruning improve things but often lost semantics and coherence with dialogue relevance. Our proposed Dynamic Neurons Masking stands out by consistently bringing the bias scores down across all languages in PCT and BBQ(singleTurn) and across all task types in FairMT-MultiTurn. Our framework not only reduces bias more effectively but also works reliably in complex, real dialogue situations. In the baseline, the effective bias score rises steadily with each turn, indicating that bias accumulates as the conversation continues.
Whereas, Table 2 demonstrates consistent bias reduction across all evaluated languages, with the largest improvements in challenging cases such as Sindhi and Pashto, resulting in more stable and reliable responses across multilingual settings. In contrast, the mitigated curve grows more slowly, demonstrating that masking effectively dampens bias propagation across turns. The shaded region represents the extent of bias reduction achieved by masking, clearly showing that the method consistently lowers bias throughout the dialogue. This highlights the effectiveness of masking in reducing bias carry-over while maintaining dialogue progression as shown in Figure 5. This illustrates how the model learned and remembered and how the knowledge remained intact; the model remembers knowledge that evolves over successive dialogue turns. Starting from a relatively low value at the first turn, the score rises steadily with each additional turn, showing that the model retains more information from previous context as the conversation progresses. The shaded confidence interval indicates that while there is some variation across conversations, the overall trend is consistent: memory carry-over strengthens over time. In Figure 6, the pattern highlights both the benefit of improved contextual consistency and the potential drawback of increased bias persistence in longer dialogues.